Which of the following values represent the definite integral 

"\intop\begin{matrix}
 1 \\
 0 
\end{matrix}" "sqrt(1- x^2) dx"


1) "\pi/2"

2) "\pi/3"

3) "\pi/3"

4) "3\pi/4"



1
Expert's answer
2021-08-23T16:05:46-0400

"\int^1_0 \sqrt{1-x^2}dx\\=[\frac{x}{2}\sqrt{1-x^2}+\frac{1}{2}sin^{-1}(x)]^1_0\\
=\frac{\pi}{4}"

reference:

1)

https://www.stumblingrobot.com/2015/10/22/use-integration-by-parts-to-establish-formulas-for-the-integral-of-the-square-root-of-1-x2/


2)

https://www.emathzone.com/tutorials/calculus/integral-of-1-over-square-root-of-1-x2.html